OpenStudy (anonymous):

Reuben runs 4 miles in 30 minutes. At the same rate, how many miles would he run in 81 minutes?

jimthompson5910 (jim_thompson5910):

solve for x 4/30 = x/81

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

I honestly don't know what would the next step be?

jimthompson5910 (jim_thompson5910):

cross multiply to get this 4*81 = 30x

jimthompson5910 (jim_thompson5910):

does that help?

jimthompson5910 (jim_thompson5910):

if not, then you can go further to get 324 = 30x 30x = 324 x = ???

OpenStudy (anonymous):

10.8

jimthompson5910 (jim_thompson5910):

getting the same thing

OpenStudy (anonymous):

okay thank you :)

jimthompson5910 (jim_thompson5910):

yw